OpenStudy (anonymous):

Julia pays a flat rate of $106 for her cell phone and is charged $0.12 for every text she sends. Julia spends at least $142 on her phone bill each month. Which of the following describes the number of texts she sends? a minimum of 300 a maximum of300 more than 300 fewer than 300

OpenStudy (anonymous):

@crush_299 this still doesnt make sense?

OpenStudy (anonymous):

@mebs could you help

OpenStudy (anonymous):

She has a flat rate of $106 that means she always pays that amount. For every text she sends she gets charged 12 cents. She spends at least 142 dollars make an equation let x = number of texts 106 + 0.12x > or equal to 142 solve the inequality. \[0.12x \ge 142 - 106\]

OpenStudy (anonymous):

so what is x 's solution from the inequality?

OpenStudy (anonymous):

i really dont know sorry

OpenStudy (anonymous):

300

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

So what does the symbol \[\ge \] mean? Yes your right

OpenStudy (anonymous):

greater or equal to

OpenStudy (anonymous):

than*

OpenStudy (anonymous):

so her text messages send are greater than or equal to 300... Can you tell me that in plain English?

OpenStudy (anonymous):

a minimum of 300

OpenStudy (anonymous):

good your done!

OpenStudy (anonymous):

thanks :) @mebs

OpenStudy (anonymous):

=)) yw

OpenStudy (anonymous):

could u help me with this one? @mebs A group of 11 people are going to the movies one Saturday. Adult tickets cost $11.50 each and child tickets cost $7.25 each. If the total cost for admission to the movies is $105.25, how many children attended the movie? 4 5 6 7

OpenStudy (anonymous):

okay so a = # of adults and c = # of children a + c = 11 and 11.50a + 7.25c = 105.25 @mebs

OpenStudy (anonymous):

im stuck between B and C @mebs

OpenStudy (anonymous):

Next time post it as a new question. c = 5 a = 6